OpenStudy (anonymous):

Solve for n. 6 – 10n = -1 – 8n – 9n

OpenStudy (mollynicole2015):

Simplifying 6 + -10n = -1 + -8n + -9n Combine like terms: -8n + -9n = -17n 6 + -10n = -1 + -17n Solving 6 + -10n = -1 + -17n Solving for variable 'n'. Move all terms containing n to the left, all other terms to the right. Add '17n' to each side of the equation. 6 + -10n + 17n = -1 + -17n + 17n Combine like terms: -10n + 17n = 7n 6 + 7n = -1 + -17n + 17n Combine like terms: -17n + 17n = 0 6 + 7n = -1 + 0 6 + 7n = -1 Add '-6' to each side of the equation. 6 + -6 + 7n = -1 + -6 Combine like terms: 6 + -6 = 0 0 + 7n = -1 + -6 7n = -1 + -6 Combine like terms: -1 + -6 = -7 7n = -7 Divide each side by '7'. n = -1 Simplifying n = -1 We good so far?

OpenStudy (anonymous):

Um, yes. Yes, I got it so far.

OpenStudy (mollynicole2015):

Now do what we did last time: Prove that it's the right answer by what?

OpenStudy (anonymous):

Putting the answer back into the equation.

OpenStudy (mollynicole2015):

So do that.

OpenStudy (anonymous):

6 – 10 x -1 = -1 – 8 x -1 – 9 x -1 6 - 10 x -1 = 16 -1 - 8 x -1 - 9 x -1 = 16 16 = 16 So our equation is true.

OpenStudy (mollynicole2015):

Great job! ;)

OpenStudy (anonymous):

Thanks for the help. :)

OpenStudy (mollynicole2015):

My pleasure!